OpenStudy (anonymous):

How to solve x^ - 18x + 80 = 0 using factoring ????

OpenStudy (anonymous):

you can see that 80=10*8 and 18=10+8 so the facto will look like (x +- 10)(x +-8) now you just need to work out if it is a + or a - it is -18x so has to be -10-8 (also -10*-8=80) so the solution (x-10)(x-8)=0

OpenStudy (anonymous):

okay thank you